# TokenChain: A Discrete Speech Chain via Semantic Token Modeling

### v1

@@ 1 Introduction
Human speech is a bidirectional mapping between symbolic text and acoustic realizations; coupling perception and production improves learning. The machine speech chain operationalizes this by training automatic speech recognition (ASR) and text-to-speech (TTS) in a closed loop [1]. Prior work enabled backpropagation from TTS to ASR via straight-through (ST) estimators [2] and achieved domain adaptation [3], underscoring this paradigm’s promise. These systems largely employed continuous intermediates (mel-spectrograms, waveforms), presenting an opportunity to harness the shift toward token-centric sequence modeling.
@@
For ASR, scaling in self-supervised learning have driven gains, and discretizing SSL features emerged as viable inputs. Early unit-based systems required language models to match log-mel baselines [12]; recent work quantizes SSL features into compact codebooks to achieve competitive WER with lower storage and I/O [13]. These results suggest semantic token sequences are suitable carriers for recognition and, by symmetry, promising targets for text-conditioned generation.
We revisit the speech chain in a fully discrete setting that (i) leverages mature discrete ASR/TTS components, (ii) preserves end-to-end feedback through discrete estimation, (iii) prioritizes semantic learning over costly acoustic synthesis.
TokenChain (Fig. 1) employs SOTA-competitive architectures, coupling a semantic-token ASR with a two-stage TTS: an autoregressive text-to-semantic (T2S) model co-trained with ASR, and a masked-generative semantic-to-acoustic (S2A) module for synthesis only. The ASR–T2S interface is textual, while feedback is entirely token-based: ST-argmax/Gumbel–Softmax enable backpropagation, and a semantic-token reconstruction loss is dynamically balanced with ASR cross-entropy. In chained training, TokenChain outperforms baselines, converging 2–6 epochs earlier with 5–13% lower equal-epoch error, and under domain adaptation reduces WER by 56% for ASR and 31% for T2S with minimal forgetting, demonstrating that the speech-chain paradigm reinstantiates effectively in the discrete-token era.
@@ 3.3 Main Results on LibriSpeech
Table 1 presents equal-compute CER/WER at epoch 12, when TokenChain variants have effectively converged and exceed baseline’s final figures at epoch 20. Across all splits, chain feedback outperforms the $L_{\text{ASR}}$-only baseline, adding to the baseline’s roughly halved error rate from the pre-chain checkpoint. The strongest variant, ST-Gumbel Anneal, yields further 10–13% relative gains on clean sets and 5–9% on other sets. Fixed $\tau$ around 1.5 is competitive, whereas sharper $\tau{\leq}1.0$ are less effective yet still surpass the baseline.
@@
Table 2 evaluates T2S by calculating Whisper-WER, SIM-O, and predicted MOS on audio synthesized with a fixed S2A. Chain feedback improves content robustness while preserving perceptual quality. ST-argmax achieves the lowest Whisper-WER (-11.6% vs. baseline). SIM-O remains within $\pm$0.5 of baseline, and MOS is stable, peaking at $\tau{=}1.5$ (+0.06). Sharper temperatures ($\tau{\leq}1.0$) degrade WER with minimal change in SIM-O/MOS, indicating that over-discrete interfaces (small $\tau$) impede text controllability of the T2S.
@@ 3.4 Domain Adaptation on TED-LIUM
Table 3 reports final-epoch CER/WER, as the dominant TED-LIUM effect is improved generalization rather than faster convergence. Although the $L_{\text{ASR}}$-only baseline already cuts WER by 52% from pre-chain, TokenChain provides further dev/test gains. The best configuration, ST-Gumbel at $\tau{=}0.75$, attains 6.0/6.2 CER and 12.7/12.6 WER, corresponding to 55.3% and 56.4% total reductions, respectively. Argmax and annealed Gumbel are competitive but not optimal. The trend indicates that sharper discrete interfaces (smaller $\tau$, approaching ST-argmax) tend to favor cross-domain transfer.
@@
Domain Behavior. Figure 3 quantifies the gain–loss asymmetry across domains: on TED-LIUM, correct rate gains are substantial: an increase of 7.5–7.6 on character and 16.3–16.4 on word, whereas on LibriSpeech, degradations are minimal: a decrease of 0.6–0.7 on character and 1.8–1.9 on word. This pattern indicates effective domain adaptation with limited forgetting of the source domain, and suggests that chain feedback promotes domain-invariant semantic alignment.
@@ 4 Conclusions
We introduced TokenChain, a machine speech chain with a fully discrete token interface. Using ST-argmax/Gumbel–Softmax with dynamic weight averaging, it enables end-to-end feedback between a semantic-token ASR and an AR T2S while keeping a NAR S2A fixed for synthesis. Empirically, TokenChain improves recognition under equal compute and converges 2–6 epochs earlier on LibriSpeech (5–13% lower equal-epoch error); under TED-LIUM adaptation it achieves substantial ASR (56%) and T2S WER (31%) reductions with limited forgetting. Ablations show annealed ST-Gumbel ($\tau\!:\!\,2.0{\to}0.1$) is strongest in-domain, whereas a sharper interface ($\tau{\approx}0.75$) favors cross-domain transfer. Future work includes adaptive learning rate scheduling, joint S2A training, scaling to larger multilingual corpora, and human evaluations.

### v2

@@ 1 Introduction
Human speech is a bidirectional mapping between symbolic text and acoustic realizations; coupling perception and production improves learning. The machine speech chain operationalizes this by training automatic speech recognition (ASR) and text-to-speech (TTS) in a semi-supervised closed loop [18]. Prior work enabled backpropagation from TTS to ASR via straight-through (ST) estimators [19] and achieved domain adaptation [25], underscoring this paradigm’s promise. These systems largely employed continuous intermediates (mel-spectrograms, waveforms), presenting an opportunity to harness the shift toward token-centric sequence modeling.
@@
For ASR, scaling in self-supervised learning has driven gains, and discretizing SSL features emerged as viable inputs. Early unit-based systems required language models to match log-mel baselines [2]; recent work quantizes SSL features into compact codebooks to achieve competitive WER with lower storage and I/O [6]. These results suggest semantic token sequences are suitable carriers for recognition and, by symmetry, promising targets for text-conditioned generation.
We revisit the speech chain in a fully discrete setting that (i) employs SOTA-competitive discrete ASR/TTS components, (ii) preserves end-to-end feedback through discrete estimation, (iii) reinstantiates the speech chain principle for semi-supervised closed-loop training via mutual feedback.
TokenChain (Fig. 1) couples a discrete semantic-token ASR with a two-stage TTS: an autoregressive text-to-semantic (T2S) model co-trained with ASR, and a masked-generative semantic-to-acoustic (S2A) module for synthesis only. The ASR–T2S interface is textual, while feedback is entirely token-based: ST-argmax/Gumbel–Softmax enable backpropagation, and a semantic-token reconstruction loss is dynamically balanced with ASR cross-entropy. In chained training, TokenChain outperforms baselines, converging 2–6 epochs earlier with 5–13% lower equal-epoch error, and under domain adaptation reduces WER by 56% for ASR and 31% for T2S with minimal forgetting, demonstrating that the speech-chain paradigm remains effective in the discrete-token era.
@@ 3.3 Main Results on LibriSpeech
Table 1 presents equal epoch budget (Epoch 12) CER/WER, when TokenChain variants have effectively converged and exceed baseline’s final figures at Epoch 20. Across all splits, chain feedback outperforms the $L_{\text{ASR}}$-only baseline, adding to the baseline’s roughly halved error rate from the pre-chain checkpoint. The strongest variant, ST-Gumbel Anneal, yields further 10–13% relative gains on clean sets and 5–9% on other sets. Fixed $\tau$ around 1.5 is competitive, whereas sharper $\tau{\leq}1.0$ are less effective yet still surpass the baseline.
@@
Table 2 evaluates T2S by calculating Whisper-WER, SIM-O, and predicted MOS on audio synthesized with a fixed S2A. Chain feedback with an apt $\tau$ can improve content robustness while preserving perceptual quality. ST-argmax achieves the lowest Whisper-WER (-11.6% vs. baseline). SIM-O remains within $\pm$0.5 of baseline, and MOS is stable, peaking at $\tau{=}1.5$ (+0.06). Sharpening ST-Gumbel ($\tau{\leq}1.0$) degrades WER with minimal change in SIM-O/MOS, indicating that sharp Gumbel interfaces (small $\tau$) impede text controllability of the T2S.
@@ 3.4 Domain Adaptation on TED-LIUM
Table 3 reports final-epoch CER/WER, as the dominant TED-LIUM effect is improved generalization rather than earlier error-rate attainment. Although $L_{\text{ASR}}$-only baseline already cuts WER by 52% from pre-chain, TokenChain provides further dev/test gains. The best setting, ST-Gumbel at $\tau{=}0.75$, attains 6.0/6.2 CER and 12.7/12.6 WER, corresponding to 55.3% and 56.4% total reductions, respectively. Argmax and annealed Gumbel are competitive but not optimal. The trend thus indicates that within ST-Gumbel, moderately sharper interfaces (smaller $\tau$) tend to facilitate cross-domain transfer.
@@
Domain Behavior. Figure 3 quantifies the gain–loss asymmetry across domains: on TED-LIUM, correct rate gains are substantial: an increase of 7.5–7.6 on character and 16.3–16.4 on word, whereas on LibriSpeech, degradations are minimal: a decrease of 0.6–0.7 on character and 1.8–1.9 on word. This pattern indicates effective domain adaptation with limited forgetting of the source domain, showing that closed-loop chain feedback promotes domain-invariant semantic alignment.
@@ 4 Conclusions
We introduced TokenChain, a discrete semantic-token closed-loop machine speech chain. Using ST-argmax/Gumbel–Softmax with dynamic weight averaging, it enables end-to-end feedback between a semantic-token ASR and an AR T2S while keeping a NAR S2A fixed for synthesis. Empirically, TokenChain improves recognition under equal epoch budget, converging 2–6 epochs earlier on LibriSpeech (5–13% lower equal-epoch error); under TED-LIUM adaptation it achieves substantial ASR (56%) and T2S WER (31%) reductions with limited forgetting. Ablations show annealed ST-Gumbel ($\tau\!:\!\,2.0{\to}0.1$) is strongest in-domain, whereas a sharper interface ($\tau{\approx}0.75$) favors cross-domain transfer. Future work includes adaptive learning rate scheduling, joint S2A training, scaling to larger multilingual corpora, and human evaluations.